OpenStudy (anonymous):

A punter kicks a football upward with an initial speed of 48 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt - 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

they gave you r = 48 so you now have h = 48t - 16t^2 ground level would be a height of 0 so h = 0 so solve for 0 = 48t - 16t^2

OpenStudy (sleepyhead314):

hint: factor out the (-16t)

OpenStudy (anonymous):

i facctored out 16 and got 0=16(3-t^2) but i dont think this is even an answer?

OpenStudy (sleepyhead314):

did ya see my hint? :P it would help to take out one of the " t "s too

OpenStudy (anonymous):

0=-16t(-3+t) or 0=16t(3-t) ?

OpenStudy (anonymous):

i dont know what to do after this though?

OpenStudy (sleepyhead314):

either or would do fine :) you should be able to get the answer out of them anyways well, for example how would you solve something like (x+1)(x-3) = 0 ? it would be kind of the same idea :)

OpenStudy (anonymous):

it has to be a numerical answer...

OpenStudy (sleepyhead314):

I know, you'll get a numerical number do you know how to get an answer out of something like (x+2)(x-3)= 0?

OpenStudy (anonymous):

x=-2 and x=3. so the answer is 3 because theres no such thing as -2 seconds lol

OpenStudy (sleepyhead314):

mmmmmmmmmmm that was just an example (16t)(3-t)=0 same idea

OpenStudy (anonymous):

t=0 and t=3 ?

OpenStudy (sleepyhead314):

yes :) so the ball leaves the ground at time 0 and hits back on the ground at time 3

OpenStudy (anonymous):

Thank you!!!

OpenStudy (sleepyhead314):

you're welcome :)

OpenStudy (anonymous):

i know its late but i have just one more and i promise i'll let you go to bed...? >.<

OpenStudy (sleepyhead314):

yeah yeah >,< one more, quickly though :P

OpenStudy (anonymous):

actually i went ahead and did it! okay goodnight now!! THANK YOU SOSOSOSO MUCH FOR EVEYTHING!!!!

OpenStudy (sleepyhead314):

lol I'm glad I could help ^_^ good night <3